# BiDoRA: Bi-level Optimization-Based Weight-Decomposed Low-Rank Adaptation

### v1

@@ B.2 GPT-2
We summarize the experimental settings for the GPT-2 experiments in Table 13. The experimental configuration, particularly during the inference stage, follows the approach described by Hu et al. (2021).
Table 13: The hyperparameters we used for GPT-2 on the E2E NLG benchmark.
@@
LoRA rank ranka = ranku = $4$
LoRA $\alpha$ $32$
Label Smooth $0.1$
@@
Lower weight decay $1$
Upper weight decay $1$
Max Seq Length $512$
@@
Beam Size $10$
Length Penalty $0.9$
no repeat ngram size $4$
@@ Appendix C Baselines in Experiments
We compare BiDoRA with Full Fine-Tuning (FT), Adapter tuning (Houlsby et al., 2019), LoRA (Hu et al., 2021), and DoRA (Liu et al., 2024) in all our experiments. We provide a brief introduction to these methods here.
Full Fine-Tuning (FT) is a commonly used method for adaptation. The model is initialized with pre-trained weights and biases, and all model parameters are updated through gradient descent.
Adapter tuning (Houlsby et al., 2019) inserts layer adapters between neural modules, such as the MLP module or the self-attention module. It incorporates two fully connected layers within an adapter layer, with a nonlinearity function applied between them.
LoRA (Hu et al., 2021) adds trainable incremental update matrices to pre-trained weight matrices. Following the experimental settings of LoRA, we applied BiDoRA to $W_{q}$ and $W_{v}$ matrices (the query and value weight matrices in the self-attention module) for a fair comparison.
DoRA (Liu et al., 2024) proposes weight-decomposed adaptation, which formulates the incremental matrices as a product of magnitude and direction components, thereby accelerating training and aligning the training behavior with full fine-tuning. In contrast, our BiDoRA trains the two components on distinct sub-datasets to alleviate overfitting.
Appendix D Theoretical Insights
We can gain several insights into how BiDoRA can alleviate the overfitting issue by analyzing both the gradient of the bi-level optimization and theoretical results from existing work.
In terms of gradient analysis, we can demonstrate the robustness of our method by examining the behavior of the loss $\mathcal{L}_{val}^{\prime}(\mathcal{V},\mathcal{M})=\mathcal{L}_{val}(\mathcal{V}-\xi\nabla_{\mathcal{V}}{\mathcal{L}_{tr}(\mathcal{V},\mathcal{M})},\mathcal{M})$. We analyze its gradient with respect to the magnitude parameters $\mathcal{M}$ using the chain rule,
$\frac{d\mathcal{L}_{val}^{\prime}}{d\mathcal{M}}=\frac{\partial\mathcal{L}_{val}^{\prime}}{\partial\mathcal{M}}+\frac{\partial\bar{\mathcal{V}}}{\partial\mathcal{M}}\times\frac{\partial\mathcal{L}_{val}^{\prime}}{\partial\bar{\mathcal{V}}}$ (9)
The second term on the right-hand side, particularly the term $\frac{\partial\bar{\mathcal{V}}}{\partial\mathcal{M}}$, referred to as the best-response Jacobian in the literature, captures how the direction component would react to changes in the magnitude. More specifically, the update of $\mathcal{M}$ must consider not only the current value of the direction $\mathcal{V}$ but also additional information about how the direction responds to changes in the magnitude. This facilitates finding a globally optimal magnitude value, thereby enhancing stability and robustness. Consequently, the model becomes more resilient to overfitting. In contrast, if $\xi$ is set to 00, Algorithm 1 reduces to a first-order approximation, which performs coordinate descent for $\mathcal{M}$ and $\mathcal{V}$ on different empirical distributions. In this scenario, the second term in Eq. (9) vanishes, making the approach insufficient to effectively alleviate overfitting.
On the other hand, we can also see the generalization guarantees of BiDoRA using various theoretical works in the literature (Oymak et al., 2021; Bai et al., 2021; Bao et al., 2021; Chen et al., 2022; Huang et al., 2022). For example, Theorem 1 in Oymak et al. (2021) provides strong evidence of the benefits of applying bi-level optimization to alleviate overfitting. Oymak et al. (2021) discussed the generalization guarantees for neural architecture search with a train-validation split. They assumed the same scenario as ours: using bi-level optimization where one optimizes the weights over the training data (lower-level problem) and hyperparameters over the validation data (upper-level problem). The theorem shows that as soon as the size of the validation data exceeds the effective number of hyperparameters (up to logarithmic factors), the test error is close to the validation error (i.e., validation error is indicative of the test error). In practice, the number of magnitude parameters is generally less than or comparable to the size of the validation set, making this assumption feasible. We empirically show that validation performance and test performance align, as theoretically supported by Oymak et al. (2021). In Figure 4, we plot the performance on the training, validation, and test datasets for the CoLA and SST-2 datasets during the search phase. The first observation is that training performance always overfits until reaching zero error, whereas validation performance closely tracks test performance. This training behavior aligns with the fact that large-capacity networks can perfectly fit (Du et al., 2019; Ji & Telgarsky, 2019; Oymak & Soltanolkotabi, 2020) in the overparameterized setting, where the available data (less than $10$k for CoLA and $67$k for SST-2) is less than the number of learnable parameters (around $170$k). To truly find a model that achieves good generalization performance, the optimization procedure should evaluate the generalization loss. Thus, as shown in Figure 4, the validation phase serves as a crucial test proxy in the overparameterized setting where training performance may not be indicative of test performance.
Figure 4: Train/Validation/Test performance during the search phase of fine-tuning the RoBERTa-base model on the CoLA dataset (left) and SST-2 (right).
Appendix E Weight Decomposition Analysis
@@
We provide a brief review of the weight decomposition analysis proposed in Liu et al. (2024). Define the weight decomposition of a weight matrix $W\in\mathbb{R}^{d\times k}$ (e.g., query matrix in an attention layer) as $W=m\frac{V}{\|V\|_{c}}=\|W\|_{c}\frac{W}{\|W\|_{c}}$, where $m\in\mathbb{R}^{1\times k}$ is the magnitude vector, and $V\in\mathbb{R}^{d\times k}$ is the directional matrix, with $\|\cdot\|_{c}$ representing the vector-wise norm of a matrix across each column. This decomposition ensures that each column of $V/\|V\|_{c}$ remains a unit vector, and the corresponding scalar in $m$ defines the magnitude of each vector. Liu et al. (2024) examine the magnitude and directional variations between $W_{0}$ and $W_{\mathrm{FT}}$, defined as $\Delta M_{\mathrm{FT}}^{t}=\frac{\sum_{n=1}^{k}|m_{\mathrm{FT}}^{n,t}-m_{0}^{n}|}{k}$ and $\Delta D_{\mathrm{FT}}^{t}=\frac{\sum_{n=1}^{k}(1-\cos(V_{\mathrm{FT}}^{n,t},W_{0}^{n}))}{k}$. Here, $\Delta M_{\mathrm{FT}}^{t}$ and $\Delta D_{\mathrm{FT}}^{t}$ represent the magnitude and direction differences between $W_{0}$ and $W_{\mathrm{FT}}$ at the $t$-th training step, respectively, with $\cos(\cdot,\cdot)$ denoting cosine similarity. $m_{\mathrm{FT}}^{n,t}$ and $m_{0}^{n}$ are the $n^{th}$ scalars in their respective magnitude vectors, while $V_{\mathrm{FT}}^{n,t}$ and $W_{0}^{n}$ are the $n^{th}$ columns in $V_{\mathrm{FT}}^{t}$ and $W_{0}$. Intuitively, a consistent positive slope trend across all the intermediate steps implies a difficulty in concurrent learning of both magnitude and direction, suggesting that slight directional changes are challenging to execute alongside more significant magnitude alterations. In contrast, a relatively negative slope signifies a more varied learning pattern, with a more pronounced negative correlation indicating a larger learning capacity.
@@
Complementary to Figure 3 in the main paper on the query matrix, we provide additional results of weight decomposition analysis in Figure 5 on the value matrix to complement the findings in Section 5.6. We can draw two key observations from Figure 5: 1) Consistent with the results in Liu et al. (2024), both FT and DoRA exhibit negative correlation values of $-49.279$ and $-5.485$, respectively, while LoRA shows a positive correlation with a value of $5032$) BiDoRA achieves a negative correlation value of $-10.547$, indicating closer alignment with FT compared to DoRA. The analysis of how BiDoRA achieves this improvement is similar to that discussed in Section 5.6.
(a) FT($k=-49.279$)
(b) LoRA($k=2.503$)
(c) DoRA($k=-5.485$)
(d) BiDoRA($k=-10.5$)
@@
Figure 5: Magnitude and direction updates for (a) FT, (b) LoRA, (c) DoRA, and (d) BiDoRA of the value matrices across different layers and intermediate steps after fine-tuning the GPT2 model on the E2E dataset. Different markers represent matrices from different training steps, while different colors indicate matrices from each layer. The values of negative correlation are shown at the top, denoted by $k$.
@@ Appendix F The Role of Hyperparameter
The hyperparameter tuning for BiDoRA is simple, convenient, and straightforward. We further conducted experiments regarding the dataset partition of $\mathcal{D}_{tr}$ and $\mathcal{D}_{val}$ to provide insights into its role in BiDoRA.
The dataset partition helps maintain the balance of inner/outer optimization by assigning different portions of data. The direction component has more trainable parameters, so it is reasonable to use more data for training the lower level while using the remaining data for training magnitudes. As shown in Table 14, We varied the inner-level dataset $\mathcal{D}_{tr}$ partition from $0.6$ to $1.0$ with $0.1$ intervals and experimented with RoBERTa-base on three splits of the Reuters21578 dataset to examine its influence.
The results indicate that both extreme cases are negative to the overall performance. When the inner partition is too small ($\leq 0.6$), directions are not well-trained, and when the inner partition is $1.0$, magnitudes are not trained at all, leading to a significant performance drop. These findings demonstrate that bi-level optimization is effective in the sense that both levels are necessary for enhancing performance. Although tuning the partition ratio may further improve overall performance, we maintain a consistent data partition of $8$:$2$ in all the experiments for simplicity. A fixed configuration of data partition already consistently yields superior performance of BiDoRA, demonstrating that our method is robust to this hyperparameter within a certain range.
@@
$0.6$ $85.32$ $79.76$ $77.69$
$0.7$ $85.32$ 80.01 77.74
$0.8$ 85.34 $79.93$ $77.63$
$0.9$ $85.27$ $79.85$ $77.64$
$1.0$ $85.23$ $79.59$ $77.42$
Table 14: Experiment results on different data partitions of BiDoRA

### v2

@@ B.2 GPT-2
We summarize the experimental settings for the GPT-2 experiments (Table˜12) in Table˜8. The experimental configuration, particularly during the inference stage, follows the approach described by Hu et al. (2022b).
Table 8: The hyperparameters we used for GPT-2 on the E2E NLG benchmark (Novikova et al., 2017).
@@
LoRA rank ranka = ranku = 444
LoRA $\alpha$ 323232
Label Smooth 0.10.10.1
@@
Lower weight decay 111
Upper weight decay 111
Max Seq Length 512512512
@@
Beam Size 101010
Length Penalty 0.90.90.9
no repeat ngram size 444
@@ Appendix C Baselines in Experiments
Here, we provide a brief introduction to compare baselines in all our experiments.
  • Full Fine-Tuning (FT): The entire model is fine-tuned, with updates to all parameters.
  • Adapter Tuning (Houlsby et al., 2019; Lin et al., 2020; Rücklé et al., 2021; Pfeiffer et al., 2021): Methods that introduce adapter layers between the self-attention and MLP modules for parameter-efficient tuning.
  • LoRA (Hu et al., 2022a): A method that estimates weight updates via low-rank matrices.
  • AdaLoRA (Zhang et al., 2023): An extension of LoRA that dynamically reallocates the parameter budget based on importance scores.
  • DoRA (Liu et al., 2024a): Decomposes pretrained weights into magnitude and direction, using LoRA for efficient directional updates.
  • VeRA (Kopiczko et al., 2024): Employs a single pair of low-rank matrices across all layers to reduce trainable parameters.
  • FourierFT (Gao et al., 2024): Fine-tunes models by learning a subset of spectral coefficients in the Fourier domain.
  • AFLoRA (Liu et al., 2024b): Freezes low-rank adaptation parameters using a learned score, reducing trainable parameters while maintaining performance.
  • LaMDA (Azizi et al., 2024): Fine-tunes large models via spectrally decomposed low-dimensional adaptation.
  • SSH (Shen et al., 2025b): Fine-tunes large models after transforming weight matrices with the discrete Hartley transformation (DHT).
  • MaCP (Shen et al., 2025a): Fine-tunes large models by projecting the low-rank adaptation weight change into the discrete cosine space.
Appendix D Weight Decomposition Analysis
@@
We provide a brief review of the weight decomposition analysis proposed in Liu et al. (2024a). Define the weight decomposition of a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times k}$ (e.g., query matrix in an attention layer) as $\mathbf{W}=\mathbf{m}\frac{\mathbf{V}}{\|\mathbf{V}\|_{c}}=\|\mathbf{W}\|_{c}\frac{\mathbf{W}}{\|\mathbf{W}\|_{c}}$, where $\mathbf{m}\in\mathbb{R}^{1\times k}$ is the magnitude vector, and $\mathbf{V}\in\mathbb{R}^{d\times k}$ is the directional matrix, with $\|\cdot\|_{c}$ representing the vector-wise norm of a matrix across each column. This decomposition ensures that each column of $\mathbf{V}/\|\mathbf{V}\|_{c}$ remains a unit vector, and the corresponding scalar in $\mathbf{m}$ defines the magnitude of each vector. Liu et al. (2024a) examine the magnitude and directional variations between $\mathbf{W_{0}}$ and $\mathbf{W}_{\mathrm{FT}}$, defined as $\Delta\mathbf{M}_{\mathrm{FT}}^{t}=\frac{\sum_{n=1}^{k}|\mathbf{m}_{\mathrm{FT}}^{n,t}-\mathbf{m}_{0}^{n}|}{k}$ and $\Delta\mathbf{D}_{\mathrm{FT}}^{t}=\frac{\sum_{n=1}^{k}(1-\cos(\mathbf{V}_{\mathrm{FT}}^{n,t},\mathbf{W_{0}}^{n}))}{k}$. Here, $\Delta\mathbf{M}_{\mathrm{FT}}^{t}$ and $\Delta\mathbf{D}_{\mathrm{FT}}^{t}$ represent the magnitude and direction differences between $W_{0}$ and $W_{\mathrm{FT}}$ at the $t$-th training step, respectively, with $\cos(\cdot,\cdot)$ denoting cosine similarity. $m_{\mathrm{FT}}^{n,t}$ and $m_{0}^{n}$ are the $n^{\text{th}}$ scalars in their respective magnitude vectors, while $\mathbf{V}_{\mathrm{FT}}^{n,t}$ and $\mathbf{W_{0}}^{n}$ are the $n^{\text{th}}$ columns in $\mathbf{V}_{\mathrm{FT}}^{t}$ and $\mathbf{W_{0}}$. Intuitively, a consistent positive slope trend across all the intermediate steps implies a difficulty in concurrent learning of both magnitude and direction, suggesting that slight directional changes are challenging to execute alongside more significant magnitude alterations. In contrast, a relatively negative slope signifies a more varied learning pattern, with a more pronounced negative correlation indicating a larger learning capacity.
@@
Complementary to Fig.˜3 in the main paper on the query matrix, we provide additional results of weight decomposition analysis in Fig.˜4 on the value matrix to complement the findings in Section˜5.5. We can draw two key observations from Fig.˜4: 1) Consistent with the results in Liu et al. (2024a), both FT and DoRA exhibit negative correlation values of 49.279-49.279- 49.279 and 5.485-5.485- 5.485, respectively, while LoRA shows a positive correlation with a value of 2.5032.5032.503. 2) BiDoRA achieves a negative correlation value of 10.547-10.547- 10.547, indicating closer alignment with FT compared to DoRA. The analysis of how BiDoRA achieves this improvement is similar to that discussed in Section˜5.5.
[…]
@@
Figure 4: Magnitude and direction updates for (a) FT, (b) LoRA, (c) DoRA, and (d) BiDoRA of the value matrices across different layers and intermediate steps after fine-tuning the GPT2 model on the E2E dataset (Novikova et al., 2017). Different markers represent matrices from different training steps, while different colors indicate matrices from each layer. The values of negative correlation are shown at the top, denoted by $k$.
Appendix E Training Cost
Table 9: Average training time cost on the GLUE benchmark (Wang et al., 2019).
Method LoRA DoRA BiDoRA
Per-step cost ×1$\times$ 1× 1 ×1.36$\times$ 1.36× 1.36 ×3.54$\times$ 3.54× 3.54
Total steps 27.4527.4527.45k 27.4527.4527.45k 17.3717.3717.37k
Total time ×1$\times$ 1× 1 ×1.36$\times$ 1.36× 1.36 ×2.24$\times$ 2.24× 2.24
Table˜9 compares the training efficiency of LoRA, DoRA, and BiDoRA on the GLUE benchmark using the RoBERTa-base model. The table details the total training steps required for convergence and the per-step computational cost, which is normalized relative to LoRA for reference. For a fair comparison, all methods were benchmarked on a single NVIDIA A100 GPU. The results show that BiDoRA converges in fewer steps than LoRA and DoRA, while the per-step cost for BiDoRA is modestly higher, as its BLO process requires iterative updates between the two levels and the computation of hypergradients. The total training time for BiDoRA is approximately 1.641.641.64 times that of DoRA, a training cost that remains comparable to the baselines. Given BiDoRA’s superior performance across various tasks, we argue that this slight increase in computational cost is an acceptable trade-off, underscoring our method’s practicality.
@@ Appendix F The Role of Hyperparameter
The hyperparameter tuning for BiDoRA is simple, convenient, and straightforward. We further conducted experiments regarding the dataset partition of $\mathcal{D}_{tr}$ and $\mathcal{D}_{val}$ to provide insights into its role in BiDoRA. The dataset partition helps maintain the balance of inner/outer optimization by assigning different portions of data. The direction component has more trainable parameters, so it is reasonable to use more data for training the lower level while using the remaining data for training magnitudes. As shown in Table˜10, we varied the inner-level dataset $\mathcal{D}_{tr}$ partition from 0.60.60.6 to 1.01.01.0 with 0.10.10.1 intervals and experimented with RoBERTa-base on three splits of the Reuters21578 dataset to examine its influence.
The results indicate that both extreme cases are negative to the overall performance. When the inner partition is too small (0.6$\leq$ 0.6≤ 0.6), directions are not well-trained, and when the inner partition is 1.01.01.0, magnitudes are not trained at all, leading to a significant performance drop. These findings demonstrate that BLO is effective in the sense that both levels are necessary for enhancing performance. Although tuning the partition ratio may further improve overall performance, we maintain a consistent data partition of 8:28:28 : 2 in all the experiments for simplicity. A fixed configuration of data partition already consistently yields superior performance of BiDoRA, demonstrating that our method is robust to this hyperparameter within a certain range.
Table 10: Experiment results on different data partitions of BiDoRA.
@@
0.60.60.6 85.3285.3285.32 79.7679.7679.76 77.6977.6977.69
0.70.70.7 85.3285.3285.32 80.01 77.74
0.80.80.8 85.34 79.9379.9379.93 77.6377.6377.63
0.90.90.9 85.2785.2785.27 79.8579.8579.85 77.6477.6477.64
1.01.01.0 85.2385.2385.23 79.5979.5979.59 77.4277.4277.42
Appendix G Comparison with Other General Methods for Addressing Overfitting
There are some common experimental settings that may be used to reduce overfitting. For DoRA, two promising methods are increasing weight decay and adopting a more aggressive dropout rate. We conducted experiments on these two methods separately. We kept hyperparameters that have been well-tuned in DoRA and can achieve optimal results while only tuning the weight decay value. Similarly, we tune the dropout rate of DoRA while keeping the weight decay value to be optimized. We conducted experiments on RoBERTa-base on three datasets. The results are presented in Table˜11.
We can draw the observation that neither of these approaches effectively addresses overfitting issues or enhances the model’s generalization ability. On the other hand, BiDoRA exploits the specific magnitude-direction structure of DoRA and the strategy of training the two distinct components on separate splits of the dataset. An advantage of our methodology is that it can be easily combined with other general-purpose overfitting-alleviating strategies since BiDoRA does not modify the original DoRA architecture.
Table 11: Experiment results on different weight decay values and different dropout rates of DoRA.
Method CoLA MRPC RTE
DoRA (weight decay = 0) 59.3 88.7 72.9
DoRA (weight decay = 0.05) 60.1 89.2 73.3
DoRA (weight decay = 0.1) 60.5 89.2 73.2
DoRA (weight decay = 0.2) 60.3 89.0 73.2
DoRA (dropout rate = 0) 59.2 89.2 72.9
DoRA (dropout rate = 0.1) 60.2 88.9 71.4
DoRA (dropout rate = 0.2) 55.1 87.8 64.2
BiDoRA 61.3 89.4 76.0
Appendix H Additional Experiments
H.1 Experiments on Natural Language Generation Tasks
In this section, we evaluate BiDoRA’s performance on the NLG task. Table˜12 presents the results of fine-tuning a GPT-2 model on the E2E dataset with baseline PEFT methods and BiDoRA. The results show that BiDoRA achieves the best performance across all five evaluation metrics, demonstrating the superiority of BiDoRA in fine-tuning pre-trained models for NLG tasks.
Table 12: Performance of BiDoRA and baseline methods for fine-tuning GPT2-medium on the E2E dataset (Novikova et al., 2017). A higher value is better for all metrics. The best results are shown in bold.
Method #Params BLEU NIST MET ROUGE-L CIDEr
FT 354.9354.9354.9M 68.068.068.0 8.618.618.61 46.146.146.1 69.069.069.0 2.382.382.38
Adapter 11.111.111.1M 67.067.067.0 8.508.508.50 45.245.245.2 66.966.966.9 2.312.312.31
LoRA 0.390.390.39M 67.167.167.1 8.548.548.54 45.745.745.7 68.068.068.0 2.332.332.33
DoRA 0.390.390.39M 67.067.067.0 8.488.488.48 45.445.445.4 70.170.170.1 2.332.332.33
BiDoRA 0.390.390.39M 69.0 8.72 46.2 70.9 2.44
H.2 Experiments on Token Classification
The effectiveness of BiDoRA can also be observed in Table˜13, which reports the results of token classification tasks. Unlike the NLU tasks discussed in the previous section, which involve classifying entire sentences and focusing on capturing global semantics, token classification requires classifying each token within a sentence, highlighting the importance of capturing local context. On the BioNLP dataset, BiDoRA consistently outperforms baseline methods by a large margin in terms of F1 score. On the CoNLL2003 dataset, BiDoRA either outperforms or matches all baseline methods across all metrics. Consistent with our previous findings, BiDoRA effectively fine-tunes pre-trained models for token classification tasks.
H.3 More Experiments on Natural Language Understanding Tasks
Table˜13 presents the results of fine-tuning RoBERTa models on the Reuters21578 datasets, a text classification task, where BiDoRA outperforms all baseline methods by an even larger margin. Notably, BiDoRA achieves performance comparable to or even better than full fine-tuning, providing further evidence of its superiority.
Table 13: RoBERTabase/large (Rb/l) with different fine-tuning methods on the Reuters21578 (Padmanabhan et al., 2016), BioNLP (Collier et al., 2004), and CoNLL2003 (Tjong Kim Sang, 2002) benchmarks. A higher value is better for all metrics. The best results are shown in bold.
Reuters21578 BioNLP CoNLL2003
Method #Params ModApte ModHayes ModLewis Accuracy Precision Recall F1 Accuracy Precision Recall F1
Rb(FT) 125.0125.0125.0M 85.485.485.4 77.677.677.6 77.177.177.1 93.993.993.9 69.069.069.0 78.978.978.9 73.673.673.6 99.399.399.3 95.795.795.7 96.396.396.3 96.096.096.0
Rb(Adapter) 0.90.90.9M 85.3 77.577.577.5 76.876.876.8 93.993.993.9 69.169.169.1 78.878.878.8 73.773.773.7 99.3 95.795.795.7 96.496.496.4 96.096.096.0
Rb(LoRA) 0.150.150.15M 84.784.784.7 74.374.374.3 74.774.774.7 93.993.993.9 69.069.069.0 78.878.878.8 73.673.673.6 99.3 95.495.495.4 96.396.396.3 95.895.895.8
Rb(DoRA) 0.170.170.17M 84.884.884.8 78.278.278.2 76.676.676.6 94.0 69.269.269.2 79.1 73.873.873.8 99.3 95.395.395.3 96.296.296.2 95.895.895.8
Rb(BiDoRA) 0.170.170.17M 85.3 79.9 77.6 93.993.993.9 71.2 78.678.678.6 74.7 99.3 95.9 96.5 96.2
Rl(FT) 355.0355.0355.0M 84.884.884.8 77.577.577.5 76.676.676.6 94.094.094.0 69.469.469.4 79.679.679.6 74.174.174.1 99.499.499.4 96.296.296.2 97.097.097.0 96.696.696.6
Rl(Adapter) 0.440.440.44M 84.884.884.8 77.977.977.9 76.776.776.7 94.0 69.469.469.4 79.779.779.7 74.274.274.2 99.4 96.196.196.1 97.097.097.0 96.696.696.6
Rl(LoRA) 0.390.390.39M 84.784.784.7 77.777.777.7 76.776.776.7 93.993.993.9 69.269.269.2 79.379.379.3 73.973.973.9 99.4 96.296.296.2 97.097.097.0 96.696.696.6
Rl(DoRA) 0.390.390.39M 84.884.884.8 77.477.477.4 76.776.776.7 94.0 69.469.469.4 79.7 74.274.274.2 99.4 96.296.296.2 97.1 96.696.696.6
Rl(BiDoRA) 0.390.390.39M 84.9 78.9 77.3 94.0 71.3 79.379.379.3 75.1 99.4 96.4 97.1 96.7
Appendix I Evidence on Orthogonality of Incremental Matrix
(a) DoRA-query matrix
(b) BiDoRA (w/o cst.)-query matrix
(c) BiDoRA-query matrix
(d) DoRA-value matrix
(e) BiDoRA (w/o cst.)-value matrix
(f) BiDoRA-value matrix
Figure 5: Eigenspectra of the direction matrix for query (top) and value (bottom) matrices across different layers. The figure compares three fine-tuning methods: BiDoRA, BiDoRA without orthogonal regularization (w/o cst.), and DoRA. Both axes are on a log scale, and only the 646464 largest eigenvalues are shown for visualization clarity. Experiments were conducted on the CoLA dataset (Warstadt et al., 2019) with the RoBERTa-base model.
To verify that the orthogonal regularization (OR) proposed in Section˜4.2 encourages the columns of the direction matrix to be orthogonal, we visualize the normalized eigenvalues of the matrix in Fig.˜5. The results show that, compared to methods without OR (i.e., DoRA and BiDoRA w/o cst.), BiDoRA with OR produces eigenvalues that are more closely aligned with those of a purely orthogonal matrix, where all eigenvalues would be one. This effect holds for both the query and value matrices and verifies the effectiveness of the OR constraint.